1. Compute nominal GDP in each year using 2017 as the base year.

2017 2018 2019

price/unit Total quantity price /unit total quantity price total quantity

good A 30 900 35 1,000 34 1,050
good B 120 192 140 200 136 205
good C 95 450 85 750 102 550


Question No 2: compute the GDP deflator in each year and Use it compute the inflation rate from 2017 to 2018, and from 2018 to 2019.

Question No 3: On the basic of data provided, Use 2018 as base year and
compute (CPI) for each year. Compare the results of inflation rates computed in Qus No 2 and 3.
1
Expert's answer
2020-10-07T07:12:47-0400
  1. Nominal GDP

2017: "(30*900)+(120*192)+(95*450)= 92790"


2018 :"(35*1000)+(140*200)+(85*750)=126750"


2019: "(34*1050)+(136*205)+(102*550)=119680"


2.GDP deflector and inflation rate

2017:"\frac{92790}{92790}*100=100\%"


2018: "\frac{126750}{92790}*100=136.599 \%"


2019: "\frac{119680}{92790}*100=128.979 \%"


inflation rate:

2018: "[136.599-100 ]\over 100" =0.36599

2019 :"[128.979-100]\over 100" =0.28979


3.computing CPI using 2018 as base year

CPI="base ~year~quantities *current~year~ price\over base~year ~basket~ quantities *base~year~price" *100

using 2018:

2017:"92790 \over 126750" =0.73207


2019:"119680 \over 126750" =0.94422


CPI is higher than inflation rate